OpenStudy (anonymous):

The distance traveled at a constant speed is directly proportional to the time of travel. If Olivia traveled 306 miles in 8.5 hours, how many miles will Olivia travel in 9.1 hours at the same constant speed? 323.6 mi 288.4 mi 327.6 mi 21.6 mi

OpenStudy (anonymous):

Cross-multiply.

OpenStudy (anonymous):

\[\frac{w}{x} = \frac{y}{z}\] \[wz = xy\] and therefore \[y=\frac{wz}{x}\]

OpenStudy (anonymous):

I don't understand which on would be x and which would be y ?

hero (hero):

General Steps: 1. Use d = rt to answer 2. input values for distance and time into the formula 3. Isolate and solve for r 4. Input second set of values for t and r into the same formula. Find d; Specific Steps 1.) d = rt 2.) 306 = r(8.5) 3.) r = 306/8.5 r = 36 4.) d = (36)(9.1) d = ?

OpenStudy (anonymous):

D= 327.6 ?

hero (hero):

Good job

OpenStudy (anonymous):

thank youuu ! (:

OpenStudy (anonymous):

@Hero's method is more versatile, so you should probably use his. But for an easier question like this, because velocity is constant, you can cross-multiply. You know she goes 306 miles in 8.5 hours, so: \[\frac{306}{8.5}\] You want to know how far she'll go in 9.1 hours: \[\frac{y}{9.1}\] \[\frac{306}{8.5} = v = \frac{y}{9.1}\] \[y = \frac{(306)(9.1)}{8.5}\] Can you solve that?

OpenStudy (anonymous):

its 327.6 (: